3) Suppose you throw a baseball straight up at a velocity of 64 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
16 represents 1/2g, the gravitational pull due to gravity (measured in feet per second2).
v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
a) What is the function that describes this problem?
Answer:



b) The ball will be how high above the ground after 1 second?
Answer:
Show work in this space.



c) How long will it take to hit the ground?
Answer:
Show work in this space.



d) What is the maximum height of the ball? What time will the maximum height be attained?
Answer:
Show work in this space.

Answer by wuwei96815(245)

a) s=-16t2+v0t

b) s=-16(1)2+(64)(1)
s=-16+64
s=48 feet

c) Ball will hit the ground when s=0.
0=-16t2+(64)t
0=-16t(t-4)
The two solutions for this equation are t=0 and t=4. Therefore, we now know that the ball will hit the ground after 4 seconds. We can ignore the t=0 as meaningless.

d) We know that the ball will reach its maximum height halfway through its flight. What is that? Half of 4 seconds is 2.
s=-16t2+v0t
s=-16(2)2+(64)(2)
s=-64+128=64 feet

I hope that I did the math correctly.